def main():
    n = int(input('Enter number of cookies: '))
    r = n / 48
    sugar = r * 1.5
    butter = r
    flour = r * 2.75
    print('You need ' + str(sugar) + ' cups of sugar, ' + str(butter) + ' cups of butter, and ' +
          str(flour) + ' cups of flour for ' + str(n) + ' cookies.')


main() 

Please explain what each function in this program does? Python
Answer #1

def main():
   #reading the number from user
n = int(input('Enter number of cookies: '))
   #dividing it 48
r = n / 48
   #finding number of sugar cups required
sugar = r * 1.5
   #finding number of butter cup required
butter = r
   #finding number of flour cup required  
flour = r * 2.75
   // displaying the all the required details for the given cookies
print('You need ' + str(sugar) + ' cups of sugar, ' + str(butter) + ' cups of butter, and ' +
str(flour) + ' cups of flour for ' + str(n) + ' cookies.')